OpenStudy (anonymous):

solution of x^2+10x+21=0

OpenStudy (saifoo.khan):

Give it a try.

OpenStudy (saifoo.khan):

we will help, where are you stuck

OpenStudy (saifoo.khan):

?

OpenStudy (anonymous):

helpp

OpenStudy (saifoo.khan):

where are you stuck

OpenStudy (anonymous):

on everything. i dont know how to do it at all

OpenStudy (anonymous):

help please?

OpenStudy (diyadiya):

x^2 + 10x +21 = 0 x^2 + 3x + 7x + 21 =0 x( x +3 ) + 7 (x + 3) =0 (x+3) (x + 7 ) =0 x+3 =0 x+7=0 x=-3 x=-7

OpenStudy (anonymous):

Rational roots\[x^3-2x^2+10x+136=0\] and thank you!

OpenStudy (anonymous):

help?